I have looked in my textbook, there wasnt anything on word problems, I tried asking my mom, my sister, and my grandmom for help and I still did not recieve any help.
I need help with finding the answer to : " 6 is 2 more than the quotient of a number and 4"

Found 2 solutions by mananth, edjones:
Answer by mananth(16946)

Let the number be x
quotient of a numberand 4 =x/4
6-2= x/4
4=x/4
multiply by 4
16=x
CHECK
6-2=4
16/4 =4
m.ananth@hotmail.ca

Answer by edjones(8007)

6 = x/4 + 2 Quotient is the result of a division of 2 numbers. 6/2=3. 3 is the quotient.
24=x+8 Multiply each side by 4.
x=16
.
Ed